It is desired to produce 2,000 kg of Fe in the reduction of its ore Fe2O3 using C as the reducing agent. What is the amount of C needed? How many Kg of the ore must be refined or reduced?

The equation: 2Fe2O3   +  3C = → 4Fe3CO

1
Expert's answer
2021-01-11T03:47:05-0500

M(Fe) = 55.84 g/mol

n(Fe) = "\frac{2000 \times 10^3}{55.84} = 35.51 \times 10^3 \;mol"

n(Fe2O3) "= \frac{1}{2}n(Fe)"

"= \frac{1}{2} \times 35.51 \times 10^3 \\

= 17.91 \times 10^3 \;mol"

M(Fe2O3) = 159.69 g/mol

m(Fe2O3) "= 17.91 \times 10^3 \times 159.69"

"= 2859.77 \times 10^3 \;g \\

= 2858.77 \;kg"

"n(C) = \frac{3}{4}n(Fe) \\

= \frac{3}{4} \times 35.51 \times 10^3 \\

= 26.63 \times 10^3 \;mol \\

m(C) = 26.63 \times 10^3 \times 12.01 \\

= 316.82 \times 10^3 \;g \\

= 316.82 \;kg"